OpenStudy (anonymous):

Integration by parts: I'm trying to identify u and dv of the following problem: 6x sinx cosx dx. (Will rewrite after the jump)

OpenStudy (anonymous):

\[\int\limits_{}^{}6x \sin(x) \cos(x) dx\] Does sin x cos x dx become dv and then from there I'll have to use the repeated method of integration by parts?

OpenStudy (asnaseer):

first use this fact to simplify the integral:\[\sin(2x)=2\sin(x)\cos(x)\]then use integration by parts.

OpenStudy (anonymous):

okay, thank you.

OpenStudy (asnaseer):

yw